# QF-tuner: Breaking Tradition in Reinforcement Learning

### v3

@@ 4 Results and Discussion
Experiments were conducted to evaluate the performance of the QF-tuner and compare it with several established optimization algorithms, including PSO, BA, and GA. OpenAI Gym environment control tasks were utilized during the experiments, as explained in section 3.1. The parameter settings for the QF-tuner comprised 100 iterations, 30 Fox agents, 10 runs, and 200 episodes, with 200 steps for the FrozenLake control task and 500 steps for the CartPole control task. The following tables 1, 2, 2 and 3 display the performance results using hyperparameters (Alpha $\alpha$, and Gamma $\gamma$), as well as Rewards, where the maximum value for FrozenLake is one and for CartPole is 500. The processing time (Time) was also considered to measure the speed of the optimization algorithm. The reward curves in the figs. 2 and 3 have been normalized between 0 and 1 to reduce the variance and improve visualization.
In fig. 2, the QF-tuner achieved a reward of 0.95, surpassing PSO (0.78), BA (0.65), GA (0.62), and Random (0.34). table 1 shows that QF-tuner reduced learning time (steps) by 33% compared to GA, 25% compared to PSO, and 20% compared to BA. Furthermore, the average increase in the reward for QF-tuner was about 36% compared to the other optimization algorithms, while the average reduction in learning time was around 26%, demonstrating its superior performance in the FrozenLake control task.
@@
Figure 2: Rewards of the QF-tuner and other optimization algorithms on the FrozenLake control task
In fig. 3, the QF-tuner achieved a reward of 499.08, surpassing PSO (398.99), GA (215.03), Random (129.46), and BA (96.65). table 2 shows that QF-tuner reduced learning time (steps) by 26% compared to GA, 20% compared to PSO, and 14% compared to BA. Moreover, the average gain in the reward for QF-tuner was approximately 57% compared to the other optimization algorithms, while the average decrease in learning time was about 20%, showing its superior performance in the CartPole control task.
@@
Figure 3: Rewards of the QF-tuner and other optimization algorithms on the CartPole control task
Additionally, QF-tuner outperformed the results of [15] by 8% in reward on the $4\times 4$ FrozenLake control task. Moreover, QF-tuner achieved a 6% higher reward and a 26% reduction in learning time compared to the study in [16] on the same task. On the other hand, QF-tuner outperformed the results of [12] by 25% in reward on the CartPole control task. Thus, these results indicate that the QF-tuner improves Q-learning performance in the FrozenLake and CartPole control tasks. The improvement in reward highlights the critical role of the proposed objective function in eq. 11 in boosting the optimization process. Furthermore, the significant reduction in learning time emphasizes the effective integration of FOX with Q-learning, leading to superior performance due to their strong compatibility. However, a key limitation of the QF-tuner is its resource-intensive nature, which prevents it from being directly applied to real-world problems. Instead, it needs to be executed in a simulated environment to fine-tune the hyperparameters, which can then be utilized directly.
@@ 5 Conclusion
In conclusion, this paper presents a novel QF-tuner method for automatic hyperparameter tuning in the Q-learning algorithm using the FOX optimization algorithm (FOX). Additionally, we propose a new objective function that combines reward, learning error, and learning time into a single function, with more emphasis on the reward. The FOX utilizes the proposed objective function to minimize the fitness value derived by executing the Q-learning algorithm. The QF-tuner has been evaluated using benchmark control tasks from the OpenAI Gym, allowing for comparisons with other methods and studies. The superior results indicate that the QF-tuner can be a feasible alternative to established hyperparameter tuning methods to solve control tasks. However, tuning hyperparameters is time-intensive and requires multiple optimization iterations to reach the best set, which restricts its use in simulation environments. This limitation presents an opportunity for future research to enhance the methodology, making it applicable to real-world applications. Furthermore, other optimizers could be utilized, or further modifications could be proposed to develop alternative variations with better performance.

### v4

@@ 4 Results and Discussion
Experiments were conducted to evaluate the performance of the QF-tuner and compare it with several established optimization algorithms, including PSO, BA, and GA. OpenAI Gym environment control tasks were utilized during the experiments, as explained in Section 3.1. The parameter settings for the QF-tuner comprised 100 iterations, 30 FA s, 10 runs, and 200 episodes, with 200 steps for the FrozenLake control task and 500 steps for the CartPole control task. The following Tables 2, 3, 2 and 3 display the performance results using hyperparameters (Alpha $\alpha$, and Gamma $\gamma$), as well as Rewards, where the maximum value for FrozenLake is one and for CartPole is 500. The processing time (Time) was also considered to measure the speed of the optimization algorithm. The reward curves in the Figures 2 and 3 have been normalized between 0 and 1 to reduce the variance and improve visualization.
In Figure 2, the QF-tuner achieved a reward of 0.95, surpassing PSO (0.78), BA (0.65), GA (0.62), and Random (0.34). Table 2 shows that QF-tuner reduced learning time (steps) by 33% compared to GA, 25% compared to PSO, and 20% compared to BA. Furthermore, the average increase in the reward for QF-tuner was about 36% compared to the other optimization algorithms, while the average reduction in learning time was around 26%, demonstrating its superior performance in the FrozenLake control task.
@@
Figure 2: Normalized rewards of the QF-tuner and other optimization algorithms on the FrozenLake task. The normalization allows for fair comparison of performance, highlighting the QF-tuner’s superior reward within fewer learning steps.
In Figure 3, the QF-tuner achieved a reward of 499.08, surpassing PSO (398.99), GA (215.03), Random (129.46), and BA (96.65). Table 3 shows that QF-tuner reduced learning time (steps) by 26% compared to GA, 20% compared to PSO, and 14% compared to BA. Moreover, the average gain in the reward for QF-tuner was approximately 57% compared to the other optimization algorithms, while the average decrease in learning time was about 20%, showing its superior performance in the CartPole control task.
@@
Figure 3: Normalized rewards of the QF-tuner and other optimization algorithms on the CartPole task. Normalization facilitates comparison, showing QF-tuner’s higher reward and reduced learning time.
Additionally, QF-tuner outperformed the results of [17] by 8% in reward on the $4\times 4$ FrozenLake control task. Moreover, QF-tuner achieved a 6% higher reward and a 26% reduction in learning time compared to the study in [18] on the same task. On the other hand, QF-tuner outperformed the results of [14] by 25% in reward on the CartPole control task. Thus, these results indicate that the QF-tuner improves Q-learning performance in the FrozenLake and CartPole control tasks. The improvement in reward highlights the critical role of the proposed objective function in Equation 11 in boosting the optimization process. Furthermore, the significant reduction in learning time emphasizes the effective integration of FOX with Q-learning, leading to superior performance due to their strong compatibility. However, a key limitation of the QF-tuner is its resource-intensive nature, which prevents it from being directly applied to real-world problems. Instead, it needs to be executed in a simulated environment to fine-tune the hyperparameters, which can then be utilized directly.
@@ 5 Conclusion
In conclusion, this paper presents the QF-tuner, a novel method for automatic hyperparameter tuning in Q-learning using the FOX optimization algorithm. A key contribution is the development of a new objective function that combines reward, learning error, and learning time with an emphasis on maximizing reward. The QF-tuner was evaluated on the OpenAI Gym control tasks, FrozenLake and CartPole. The results show that the QF-tuner outperformed established algorithms, achieving a reward of 0.95 on FrozenLake, surpassing PSO (0.78), GA (0.62), and BA (0.65), while reducing learning time by 26%. On CartPole, it achieved 499.08 reward, surpassing PSO (398.99) and GA (215.03), and reducing learning time by 20%. However, tuning hyperparameters is time-intensive and requires multiple optimization iterations to reach the best set, which restricts its use in simulation environments. Future research can improve the proposed method, making it suitable for real-world applications and investigating alternative optimizers used in the Q-learning algorithm.